OpenStudy (loser66):

Consider the equilateral triangle with vertices 1,2,3 initially placed, respectively, at \((-1/2, \sqrt3/2)\), (1,0) and \((-1/2, -\sqrt3/2)\)in the x-y plane. Give 6 2 x 2 matrices implementing all of the symmetries of this equilateral triangle. Please, help

OpenStudy (loser66):

I got just \[f_1=\left[\begin{matrix}1&0\\0&-1\end{matrix}\right]\] it reflects across x-axis (or around perpendicular bisection segment 1,3 through 2) \(f_1\) sends \(\left(\begin{matrix}1&2&3\\3&2&1\end{matrix}\right)\) I mean sends 1 to 3, 2 to 2 and 3 to 1 \[f_2=\left[\begin{matrix}-1/2&\sqrt3/2\\\sqrt3/2&-1/2\end{matrix}\right]\]counterclockwise rotates around center an angle \(2\pi/3\) \(f_2\) sends \(\left(\begin{matrix}1&2&3\\3&1&2\end{matrix}\right)\) I mean sends 1 to 3, 2 to 1 and 3 to 2 \[f_3=\left[\begin{matrix}-1/2&\sqrt3/2\\-\sqrt3/2&-1/2\end{matrix}\right]\]counterclockwise rotates around center an angle \(4\pi/3\) \(f_3\) sends \(\left(\begin{matrix}1&2&3\\2&3&1\end{matrix}\right)\) I mean sends 1 to 2, 2 to 3 and 3 to 1

OpenStudy (loser66):

\(f_4\) denotes symmetry across the perpendicular bisector from 1 to segment 23. But I don't know how to set it in matrix form, the equation of the line is \(y= \sqrt3 x\) Please, help

OpenStudy (loser66):

@SithsAndGiggles

OpenStudy (anonymous):

Sorry for taking so long, I was playing around with the Desmos calculator to see if I could graph the triangle: https://www.desmos.com/calculator/yuqwtbnzes

OpenStudy (anonymous):

Ugh these matrix transformations aren't as familiar as they should be... Is this linear algebra?

OpenStudy (loser66):

Intuitively, I can see through f4, (1,0) becomes (-1/2, sqrt3/2) :) it's not linear algebra, it's abstract algebra.

OpenStudy (loser66):

\[f_4=\left[\begin{matrix}a&b\\c&d\end{matrix}\right]\]

OpenStudy (loser66):

\[\left[\begin{matrix}a&b\\c&d\end{matrix}\right]\left[\begin{matrix}1\\0\end{matrix}\right]=\left[\begin{matrix}-1/2\\\sqrt3/2\end{matrix}\right]\]

OpenStudy (loser66):

which gives me a = -1/2 and c = sqrt3/2 but I don't know how to find b, d

OpenStudy (anonymous):

It looks like \(b\) and \(d\) can be any number (i.e. infinite number of solutions).

OpenStudy (anonymous):

Or are you supposed to get unique solutions for all constants?

OpenStudy (loser66):

Yes, I have to have a unique one.

OpenStudy (anonymous):

Is there anything preventing you from fixing \(b\) and \(d\)? We know the equation holds for any \(b,d\), so why not let them be 0 for simplicity?

OpenStudy (loser66):

It's a symmetry function, so that, it must apply for both 1 and 2. I mean 1 to 2 and 2 to 1. that means f_4 applies to (1,0) we get (-1/2, sqrt3/2) f_4 applies to (-1/2,sqrt3/2) we get (1,0) back. That's why b, d cannot be arbitrary values.

OpenStudy (anonymous):

I see, so basically, you need to satisfy \[f_n\begin{bmatrix}1\\0\end{bmatrix}=\begin{bmatrix}-1/2\\\sqrt3/2\end{bmatrix}~~\iff~~\begin{bmatrix}1\\0\end{bmatrix}={f_n}^{-1}\begin{bmatrix}-1/2\\\sqrt3/2\end{bmatrix}\] which requires \(f\) to be nonsingular, correct?

OpenStudy (loser66):

yes,

OpenStudy (loser66):

For example: the reflection about y =x will take (1,0) to (0,1) and vice versa and the matrix is \(\left[\begin{matrix}0&1\\1&0\end{matrix}\right]\) so that if applying to (1,0), we get (0,1) and IT applies to (0,1) we get (1,0)

OpenStudy (loser66):

but our line is y = sqrt3 x, ha!!

OpenStudy (anonymous):

http://planetmath.org/derivationof2dreflectionmatrix Scroll down to "Reflections across a line of given slope." In this case, \(m=\sqrt3\). I hope that's what you're looking for.

OpenStudy (loser66):

Thanks a ton. I will

OpenStudy (loser66):

@SithsAndGiggles I got it. :)

OpenStudy (anonymous):

Glad to have helped!